Neutron stars, such as the one at the center of the Crab Nebula, have about the same mass as our sun but a much smaller diameter.

If you weigh 690 N on the earth, what would be your weight on the surface of a neutron star that has the same mass as our sun and a diameter of 23.0 km ?

Take the mass of the sun to be ms = 1.99×1030 kg , the gravitational constant to be G = 6.67×10−11 N⋅m2/kg2 , and the acceleration due to gravity at the earth's surface to be g = 9.810 m/s2 .

Express your weight wstar in newtons.
Answer #1

mg = 690

m= 690/ 9.8 = 70.41kg

Acceleration due to gravity on neutron star = GMs/R^2

R = 23 x10^3/ 2= 11.5 x10^3m

g'= 6.67 x 10^-11× 1.99 x 10^30/ (11.5 x10^3)^2= 100.4 x10^10m/s^2

Water = mg'= 70.41 x 100.4 x 10^10= 7066.7 x 10^10 N